# Comfortability of quantum walks on embedded graphs on surfaces

### v1

@@ 2.1 Observation 1: In the limit of $a\to 1$
$h(x)=x\frac{1+a^{x}}{1-a^{x}}.$
Let $F=\{f_{1},\dots,f_{r}\}$, with $|f_{1}|\geq|f_{2}|\geq\cdots\geq|f_{r}|$ be the set of the underlying faces, where $r=|F|$. Then the first term of (1.1) in Theorem 1.2 can be reexpressed by
@@
Note that $|A|=|f_{1}|+\cdots+|f_{r}|$. Then the boundary lenghts of $F$ give the integer partition, which is bijective to the Young diagram. Thus in the following, let us consider what is the integer partition $\lambda\vdash|A|$ makes the first term larger. Let us the important properties of $h(x)$ for the above consideration be summarised below.
@@
For $\lambda=(x_{1},x_{2},\dots,x_{r})\vdash|A|$ with $x_{1}\geq x_{2}\geq\cdots\geq x_{r}$, we subject the condition $x_{r}\geq 3$ because the length of face is larger than $3$. For such a Young diagram $\lambda=(x_{1},\dots,x_{r})$, let us put $Q(\lambda)$ by
@@
We call it the island of the energy. Define the partially order $\lambda_{1},\lambda_{2}\vdash|A|$ by $\lambda_{1}>\lambda_{2}$ if and only if $Q(\lambda_{1})>Q(\lambda_{2})$. By property (1), large length of the Young diagram makes $Q(\cdot)$ large. By property (2), bias of sizes of rows also makes $Q(\cdot)$ large. Thus for any Young diagram $\lambda\vdash|A|$, we have
@@
Let $G$ be a connected abstract graph. If there is a triangulation in the embeddings of $G$, then that is one of the best embeddings for quantum walker.
@@
If there is a triangulation in the embeddings, then that is the most comfortable on the island. Moreover since there is no self-intersections in the triangulation, then that is the most comfortable embedding by Corollary 2.1. ∎
If $n\equiv 0,3,4,7\mod 12$, the minimum genus embedding of $K_{n}$ becomes the triangulation, because $3|F|=2|E|$ holds by Fact 1. Then we can see that Corollary 2.3 is consistent with Corollary 2.2 for $K_{n}$.
Example: The Hasse diagram in Fig. 3 describes the case for $|A|=12$ by only using the properties of $h(x)$, (1) and (2).
Figure 3: Hasse diagram with respect to the comfortability on the island.
This Hasse diagram shows the partial order of the comfortability restricted to the islands for $K_{4}$’s embeddings. In this Hasse diagraph, the order of the Young diagram $[4,4,4]$ and $[9,3]$ are not determined by only using the properties (1) and (2) of $h(x)$. However it is possible to estimate the difference between the comfortabilities on the island to $\lambda=[9,3]$ and $\mu=[4,4,4]$ in the Hasse diagram of Fig 3, $Q([9,3])$ and $Q([4,4,4])$ as follows:
$Q([9,3])<Q([5,4,3])<Q([4,4,3,1])<Q([4,4,4])+h(0).$
Here all the inequalities derive from the property (1) of $h(x)$. On the other hand,
$Q([4,4,4])<Q([5,4,3])<Q([9,3])+h(0)$
Here the first and second inequalities derive from the property (2) and (1), respectively. Then we have
@@
In the statement of Corollary 2.1, the comfortability of ``$k=2,\;[8,4]"$ is greater than that of ``$g=1,\;[8,4]"$ (see Figure 2), while these are the same by in the Hasse diagram of Fig. 3 on the island. The difference derives from the number of intersections; there are $2$ points of the intersection in the enneagon for $g=1$ while there is only $1$ point of the intersection in the enneagon for $k=2$. See Figure 4. The same reason can be applied to the case for ``$k=2,\;[9,3]"$ and ``$g=1,\;[9,3]"$. Therefore the non-orientability reduces the self-intersection and increases the comfortability.
Figure 4: The self-intersection. The comparison between the embeddings on the torus and the Klein bottle whose both faces are $1$ octagon and $1$ square. The numbers of self-intersections in the octagon at embeddings on the torus and the Klein bottle are $2$ and $1$, respectively. Thus the embedding on the Klein bottle is better than the embedding on the torus in the setting of Corollary 2.1.
@@ 3 Quick review on the rotation system
Abstract graph. Let $G(X,A)$ be a symmetric digraph with the set of the vertices $X$ and the set of the symmetric arcs $A$, that is, $e\in A$ if and only if $\bar{e}\in A$, where $\bar{e}$ is the inverse arc of $e$. In this paper, we assume there is no multiple arcs. Then an arc $e\in A$ is sometimes represented by $(o(e),t(e))$ to emphasize the origin and terminal vertices. The support edge of $e\in A$ is denoted by $|e|=|\bar{e}|$. The set of edges (which are undirected) is defined by $E=\{|e|\;|\;e\in A\}$. For each arc $e\in A$, the terminus and origin vertices are denoted by $t(e)\in X$ and $o(e)\in X$. The arc $e\in A$ with $o(e)=x$, $t(e)=y$ is also represented by $a=(x,y)\in A$. Let $A_{x}\subset A$ be the set of arcs whose terminal vertices are $x\in X$, that is, $A_{x}=\{e\in A\;|\;t(e)=x\}$.
@@
Every rotation system on graph $G$ defines (up to equivalence of embeddings) a unique locally oriented graph embedding. Conversely, every locally oriented graph embedding defines a rotation system for $G$.
The Eular’s formula gives the genus of the undelying embedding:
@@
For any closed walk $(e_{0},e_{1},\dots,e_{s-1})$ on $G$, the operation iii preserves the value of $\sum_{j=0}^{s-1}\tau(e_{j})$, which ensures that the underlying embedding is preserved by this operation iii.
@@
Figure 7: The drawing on the closed surface: Diagonally located places on the dotted boundary are identified with each other. Incident half edges of each vertex are arranged clockwise so that its rotation is conserved. Connect corresponding half-edges without any crossing and every type-00 edge passes through the dotted boundary evenly, while every type-$1$ does so oddly. In this case, there are no crossings across the dotted boundary in every the type-00, while the type-$1$ edge crosses the dotted boundary once.
@@ 4 Construction of quantum walk on the rotation system $(G,\rho,\tau)$
On the graph embedded on a closed surface, when a walker passes through a twisted edge, the rotation of the endpoint vertex is reversed. To reflect the parity of the “sheet”, that is, front/back, in the dynamics of our quantum walk, let us prepare the following notions.
@@ 4.1 Double covering
Note that the situation of a vertex $y$ with its connected edges $D_{y}=\{e_{1},\dots,e_{r}\}$ in $(G,\rho,\tau)$ is equivalent to the situation in the drawing (*) of $G^{\tau}$ that all the edges of $D_{y}$ and its chiral edges of $D_{y^{\prime}}$ with $\tau(e_{j})=0$ are in the front sheet and back sheet, respectively, while all the edges with $\tau(e_{j})=1$ cross the boundary. The operation iii. to a vertex $y$ switches the rotation and types of all its incoming and outgoing arcs. Then in the situation of $G^{\tau}$ with the drawing (*), the vertex $y$ and its chiral vertex $y^{\prime}$ are pulled to the opposite sheets conserving the locations of the other vertices. After this switching, the resulting drawing of $G^{\tau}$, the rotations of $y$ and $y^{\prime}$, and also the types of all the connected arcs are switched. ∎
@@ 4.2 Blow-up
The blow up graph of $G=(V,A)$ with the rotation $\rho$ is obtained by replacing each vertex of $G$ into the directed cycle following the rotation assigned to each vertex and conserving the adjacency relation of the original graph as follows. The vertex set of the blow up graph $G^{BU}(\rho)=(X^{BU}(\rho),A^{BU}(\rho))$ is denoted by
@@
the arc set of the blow up graph is denoted by the disjoint union of
$A^{BU}(\rho)=A_{br}\sqcup A_{is},$
which are called the bridge and the island, respectively. Here the “bridge” $A_{br}$ and the “island” $A_{is}$ are both isomorphic to $A$ and denoted by
$A_{is}=\{(e,\rho(e))\;|\;e\in A\},\;A_{br}=\{(e,\bar{e})\;|\;e\in A\}.$
See Figure 8 (c).
@@
It holds that for $e\in A$ with $t(e)=x\in X$ in $G$,
@@
We call the above directed cycle induced by $x\in X$ the island of $x\in X$ which is denoted by $A_{is}(x)$. On the other hand, for $\epsilon\in E$ with the end vertices $x,y\in X$ in $G$, $(e,\bar{e})$ and $(\bar{e},e)$, where $|e|=\epsilon$, are called the bridge between the islands $x$ and $y$. Then let us define
@@
for an island arc $\boldsymbol{e}_{is}=(e,\rho(e))$, and
$\bar{\boldsymbol{e}}_{br}:=(\bar{e},e),\;\tau(\boldsymbol{e}_{br})=\tau(|e|)$
for a bridge arc $\boldsymbol{e}_{br}=(e,\bar{e})$.
@@
The in-degree and out-degree for every vertex of the blow-up graph are equally $2$.
For each vertex $e\in X^{BU}$, the incoming arcs to $e$ come from the vertex $\rho^{-1}(e)$ along the island of $t_{G}(e)$ and the vertex $\bar{e}$ along the bridge between the islands $t(e)$ and $o(e)$ while the outgoing arcs to $e$ go out to the vertex $\rho(e)$ along the island of $t_{G}(e)$ and the vertex $\bar{e}$ along the bridge between the islands $t(e)$ and $o(e)$.
Then for any island arc $\boldsymbol{e}_{is}\in A_{is}$, there are exactly two bridge arcs $\boldsymbol{e}_{br}$ and $\boldsymbol{e}_{br}^{\prime}\in A_{br}$ such that $o(\boldsymbol{e}_{is})=t(\boldsymbol{e}_{br})$ and $t(\boldsymbol{e}_{is})=t(\boldsymbol{e}_{br}^{\prime})$, respectively. Such bridge arcs for the island arc $\boldsymbol{e}_{is}$ are denoted by
$\boldsymbol{e}_{br}:=\mathrm{br}(\boldsymbol{e}_{is}),\;\boldsymbol{e}_{br}^{\prime}:=\mathrm{br}^{\sharp}(\boldsymbol{e}_{is}),$
respectively. On the other hand, for any bridge arc $\boldsymbol{e}_{br}\in A_{br}$, there is exactly two island arcs $\boldsymbol{e}_{is}$ and $\boldsymbol{e}_{is}^{\prime}$ such that $o(\boldsymbol{e}_{br})=t(\boldsymbol{e}_{is})=o(\boldsymbol{e}_{is}^{\prime})$. Such island arcs for the bridge arc $\boldsymbol{e}_{br}$ are denoted by
$\boldsymbol{e}_{is}:=\mathrm{is}(\boldsymbol{e}_{br}),\;\boldsymbol{e}_{is}^{\prime}:=\mathrm{is}^{\sharp}(\boldsymbol{e}_{br}),$
respectively.
@@ 4.3 Quantum walk on the rotation system $(G,\rho,\tau)$
For given the rotation system $(G,\rho,\tau)$, let us consider the blow up graph of $G^{\tau}$ induced by the rotation $\rho\oplus\rho^{-1}$ by replacing $G$ discussed in the previous subsection 4.2 with $G^{\tau}$ discussed also in subsection 4.1. This blow up graph is denoted by $G(\rho,\tau)=(X(\rho,\tau),A(\rho,\tau))$. Note that $G(\rho,\tau)={G^{\tau}}^{BU}(\rho\oplus\rho^{-1})$.
@@
In our quantum walk model, a local time evolution at each vertex of $G(\rho,\tau)$ is represented by a $2$-dimensional unitary matrix, because the degree of the blow up graph is $2$, and if a quantum walk passes through the twisted bridge, the phase of the quantum walker is converted by $e^{i\pi}=-1$.
@@
with some initial state $\psi_{0}\in\mathbb{C}^{A(\rho,\tau)}$. The unitary time evolution operator $W$ is defined as follows.
@@
For any island arc $\boldsymbol{e}_{is}\in A_{is}\subset A(\rho,\tau)$, bridge arc $\boldsymbol{e}_{br}\in A_{br}\subset A(\rho,\tau)$, the time evolution operator $W$ is defined by
@@
Remark 4.3.
For any bridge arc $\boldsymbol{e}_{br}\in A_{br}\subset A(\rho,\tau)$, the local time evolution on the vertex $o(\boldsymbol{e}_{br})\in X(\rho,\tau)$ is described by

### v2

@@ 2.1 Observation 1: In the limit of $a\to 1$
$h(x)=x\frac{1+a^{x}}{1-a^{x}}$
for $10<|a|<$ 1. Here we define $h(0)$ as
$h(0)=\lim_{x\downarrow 0}h(x)=\frac{2}{\log|a|}.$
Let $F=\{f_{1},\dots,f_{r}\}$, with $|f_{1}|\geq|f_{2}|\geq\cdots\geq|f_{r}|$ be the set of underlying faces, where $r=|F|$. Then the first term of (1.1) in Theorem 1.2 can be reexpressed by
@@
Note that $|A|=|f_{1}|+\cdots+|f_{r}|$. Then the boundary lengths of $F$ provide an integer partition, which is bijective to the Young diagram. Thus in the following, we consider the integer partition $\lambda\vdash|A|$ which increases the first term. Let us summarize the important properties of $h(x)$ for the above consideration as follows. This proof can be obtained by straightforward calculations.
@@
For $\lambda=(x_{1},x_{2},\dots,x_{r})\vdash|A|$ with $x_{1}\geq x_{2}\geq\cdots\geq x_{r}$, we can impose the condition $x_{r}\geq 3$ because the length of each face is longer than $2$. For such a Young diagram $\lambda=(x_{1},\dots,x_{r})$, let us put $Q(\lambda)$ by
@@
We call this the energy of all the islands. Define the partial order $\lambda_{1},\lambda_{2}\vdash|A|$ as $\lambda_{1}>\lambda_{2}$ if and only if $Q(\lambda_{1})>Q(\lambda_{2})$. According to Property (1), a large length of the Young diagram increases $Q(\cdot)$. According to Property (2), the bias in row size also increases $Q(\cdot)$. Thus for any Young diagram $\lambda\vdash|A|$, we have
@@
Let $G$ be a connected abstract graph. If there is a triangulation in the embeddings of $G$, then it is the best embedding.
@@
If there is a triangulation in the embeddings, then it is the most comfortable on all the islands. Moreover, since there are no self-intersections in the triangulation, the triangulation is the most comfortable embedding according to Corollary 2.1. ∎
If $n\equiv 0,3,4,7\mod 12$, the minimum genus embedding of $K_{n}$ becomes the triangulation, because $3|F|=2|E|$ holds by Fact 1. Thus, we can see that Corollary 2.3 is consistent with Corollary 2.2 for $K_{n}$.
Example: The Hasse diagram in Fig. 3 describes the case for $|A|=12$ using only Properties of $h(x)$.
Figure 3: The Hasse diagram of the comfortability on the island of the embeddings.
This Hasse diagram shows the partial order of the comfortability restricted to the islands for $K_{4}$’s embeddings. In this Hasse diagram, the order of the Young diagrams $[9,3]$ and $[4,4,4]$ is not determined using only Properties (1) and (2) of $h(x)$. However it is possible to estimate the difference between $Q([9,3])$ and $Q([4,4,4])$. In fact, we have $Q[9,3]<Q[4,4]+h(0)$ by the following Hasse diagram:
$Q([9,3])<Q([5,4,3])<Q([4,4,3,1])<Q([4,4,4])+h(0),$
where all the inequalities are derived from Properties of $h(x)$ (1). On the other hand, we also have $Q[4,4]<Q[9,3]+h(0)$ by the following Hasse diagram:
$Q([4,4,4])<Q([5,4,3])<Q([9,3])+h(0),$
where the first and second inequalities are derived from Properties (2) and (1), respectively. Then we have
@@
Note that according to Fig 2 or Corollary 2.1, the comfortability of ``$k=2,\;[8,4]"$ is greater than that of ``$g=1,\;[8,4]"$, while these are the same in Fig. 3 for all the islands. The difference derives from the number of self-intersections; there are two points of the self-intersection in the enneagon for $g=1$ while there is only one point of the self-intersection in the enneagon for $k=2$. See Figure 4. The same reason can be applied to the case for ``$k=2,\;[9,3]"$ and ``$g=1,\;[9,3]"$. Therefore, the non-orientability decreases the self-intersection and increases the comfortability.
Figure 4: The self-intersection. Comparison between the embeddings on the torus and Klein bottle, whose faces are one octagon and one square. The numbers of self-intersections in the octagon at embeddings on the torus and Klein bottle are two and one, respectively. Thus the embedding on the Klein bottle is better than that on the torus in the setting of Corollary 2.1.
@@ 3 Quick review on the rotation system
Abstract graph. Let $G(X,A)$ be a symmetric digraph with the set of the vertices $X$ and the set of the symmetric arcs $A$, that is, $e\in A$ if and only if $\bar{e}\in A$, where $\bar{e}$ is the inverse arc of $e$. The support edge of $e\in A$ is denoted by $|e|=|\bar{e}|$. The set of edges (which are undirected) is defined by $E=\{|e|\;|\;e\in A\}$. In this paper, the underlying graph is assumed to be simple without self-loops. For each arc $e\in A$, the terminus and origin of $e$ are denoted by $t(e)\in X$ and $o(e)\in X$. The arc $e\in A$ with $o(e)=x$, $t(e)=y$ is also represented by $e=(x,y)\in A$. Let $A_{x}\subset A$ be the set of arcs whose terminal vertices are $x\in X$, that is, $A_{x}=\{e\in A\;|\;t(e)=x\}$.
@@
Every rotation system on graph $G$ defines (up to equivalence of embeddings) a unique locally oriented graph embedding [7]. Conversely, every locally oriented graph embedding defines a rotation system for $G$.
The Eular’s formula gives the genus of the embedding:
@@
For any closed walk $(e_{0},e_{1},\dots,e_{s-1})$ on $G$, Operation iii preserves the value of $\sum_{j=0}^{s-1}\tau(e_{j})$, which ensures that the embedding is preserved by Opeation iii.
@@
Figure 7: The drawing of $K_{4}$ on the projective plane: Diagonally located places on the dotted boundary are identified with each other. Incident half edges of each vertex are arranged clockwise so that its rotation is conserved. Connect corresponding half-edges without any crossing and every type-00 edge passes through the dotted boundary evenly, while every type-$1$ does so oddly. In this case, there are no crossings across the dotted boundary in every the type-00, while the type-$1$ edge crosses the dotted boundary once.
@@ 4 Construction of quantum walk on the rotation system $(G,\rho,\tau)$
When a walker passes through a twisted edge on the rotation system $(G,\rho,\tau)$, the rotation of the endpoint vertex is reversed. To reflect the parity of the “sheet”, that is, front/back, in the dynamics of our quantum walk, let us prepare the following notions.
@@ 4.1 Double covering
Note that the situation of a vertex $y$ with its connected edges $D_{y}=\{e_{1},\dots,e_{r}\}$ in $(G,\rho,\tau)$ is equivalent to the situation in the drawing (*) of $G^{\tau}$ that all the edges of $D_{y}$ and its chiral edges of $D_{y^{\prime}}$ with $\tau(e_{j})=0$ are in the front sheet and back sheet, respectively, while all the edges with $\tau(e_{j})=1$ cross the boundary. Operation iii to a vertex $y$ switches the rotation and types of all its incoming and outgoing arcs. Then in the situation of $G^{\tau}$ with the drawing (*), the vertex $y$ and its chiral vertex $y^{\prime}$ are pulled to the opposite sheets conserving the locations of the other vertices. After this switching, the resulting drawing of $G^{\tau}$, the rotations of $y$ and $y^{\prime}$, and also the types of all the connected arcs are switched. ∎
@@ 4.2 Blow-up
The blow-up graph of $G=(V,A)$ with the rotation $\rho$ is obtained by replacing each vertex of $G$ into the directed cycle following the rotation assigned to each vertex and conserving the adjacency relation of the original graph as follows. The vertex set of the blow-up graph $G^{BU}(\rho)=(X^{BU}(\rho),A^{BU}(\rho))$ is defined by
@@
the arc set of the blow-up graph is defined by the disjoint union of
$A^{BU}(\rho)=A_{is}\sqcup A_{br},$
which are called the bridge and the island, respectively. Here the “bridge” $A_{br}$ and the “island” $A_{is}$ are both isomorphic to $A$ and defined by
$A_{is}=\{(e,\rho(e))\;|\;e\in A\},\;A_{br}=\{(e,\bar{e})\;|\;e\in A\}$
through the direct product $A\times A$. See Figure 8 (c).
@@
It holds that, for $e\in A$ with $t(e)=x\in X$ in $G$,
@@
We call the above directed cycle induced by $x\in X$ the island of $x\in X$, which is denoted by $A_{is}(x)$. On the other hand, for $\epsilon\in E$ with the end vertices $x,y\in X$ in $G$, $(e,\bar{e})$ and $(\bar{e},e)$, where $|e|=\epsilon$, are called the bridge between the islands $x$ and $y$. Then we define an extension of $\rho$ and $\tau$ to the blow-up graph $G^{BU}(\rho)$ as follows:
@@
for an island arc $\boldsymbol{e}_{is}=(e,\rho(e))\in A_{is}$, and
$\tau(\boldsymbol{e}_{br}):=\tau(e)$
for a bridge arc $\boldsymbol{e}_{br}=(e,\bar{e})\in A_{br}$.
For each vertex $e\in X^{BU}$, the incoming arcs to $e$ come from the vertex $\rho^{-1}(e)$ along the island of $t_{G}(e)$ and the vertex $\bar{e}$ along the bridge between the islands $t_{G}(e)$ and $o_{G}(e)$ while the outgoing arcs from $e$ go out to the vertex $\rho(e)$ along the island of $t_{G}(e)$ and the vertex $\bar{e}$ along the bridge between the islands $t_{G}(e)$ and $o_{G}(e)$.
Then for any island arc $\boldsymbol{e}_{is}\in A_{is}$, there are exactly two bridge arcs incoming to $\boldsymbol{e}_{is}$: $\boldsymbol{e}_{br}$ and $\boldsymbol{e}_{br}^{\prime}\in A_{br}$ satisfy $o(\boldsymbol{e}_{is})=t(\boldsymbol{e}_{br})$ and $t(\boldsymbol{e}_{is})=t(\boldsymbol{e}_{br}^{\prime})$, respectively. Such bridge arcs for the island arc $\boldsymbol{e}_{is}$ are denoted by
$\boldsymbol{e}_{br}:=\mathrm{br}(\boldsymbol{e}_{is}),\;\boldsymbol{e}_{br}^{\prime}:=\mathrm{br}^{\sharp}(\boldsymbol{e}_{is}),$ (4.14)
respectively. On the other hand, for any bridge arc $\boldsymbol{e}_{br}\in A_{br}$, there are exactly two island arcs incident to $o(\boldsymbol{e}_{br})$: $\boldsymbol{e}_{is}$ and $\boldsymbol{e}_{is}^{\prime}$ satisfy $o(\boldsymbol{e}_{br})=t(\boldsymbol{e}_{is})=o(\boldsymbol{e}_{is}^{\prime})$. Such island arcs for the bridge arc $\boldsymbol{e}_{br}$ are denoted by
$\boldsymbol{e}_{is}:=\mathrm{is}(\boldsymbol{e}_{br}),\;\boldsymbol{e}_{is}^{\prime}:=\mathrm{is}^{\sharp}(\boldsymbol{e}_{br}),$ (4.15)
respectively. See Fig. 9.
@@
The in-degree and out-degree for every vertex of the blow-up graph are equally $2$. This condition is suitable for the purpose to implement a quantum walk model as a circuit of optical polarized elements [17].
@@ 4.3 Quantum walk on the rotation system $(G,\rho,\tau)$
For the rotation system $(G,\rho,\tau)$, let us consider the blow-up graph of $G^{\tau}$ induced by the rotation $\rho\oplus\rho^{-1}$. This blow-up graph is denoted by $G(\rho,\tau)=(X(\rho,\tau),A(\rho,\tau))$, that is,
$G(\rho,\tau):=(G^{\tau})^{BU}(\rho\oplus\rho^{-1}).$
@@
In our quantum walk model, a local time evolution at each vertex of $G(\rho,\tau)$ is represented by a $2$-dimensional unitary matrix, because the degree of the blow-up graph is $2$, and if a quantum walk passes through the twisted bridge, the phase of the quantum walker is converted by $e^{i\pi}=-1$.
@@
with some initial state $\psi_{0}\in\mathbb{C}^{A(\rho,\tau)}$. The unitary time evolution operator $W$ is defined as follows. See also Figs. 9 and 10.
@@
Let $\psi_{t}\in\mathbb{C}^{A}$ be the state of the quantum walk at time $t=0,1,2,\dots$. The time evolution operator $W$ is defined by $\psi_{t+1}=W\psi_{t}$ such that for any bridge arc $\boldsymbol{e}_{br}\in A_{br}\subset A(\rho,\tau)$, the local time evolution on the vertex $o(\boldsymbol{e}_{br})\in X(\rho,\tau)$ is described by
In other words, for any island arc $\boldsymbol{e}_{is}\in A_{is}\subset A(\rho,\tau)$ and for any bridge arc $\boldsymbol{e}_{br}\in A_{br}\subset A(\rho,\tau)$,
@@
Figure 9: The definitions of $\mathrm{is}(\boldsymbol{e}_{br})$, $\mathrm{is}^{\sharp}(\boldsymbol{e}_{br})$, $\mathrm{br}(\boldsymbol{e}_{is})$ and $\mathrm{br}^{\sharp}(\boldsymbol{e}_{is})$.
Figure 10: The local time evolution